# RapidPoseTriangulation: Multi-view Multi-person Whole-body Human Pose Triangulation in a Millisecond

### v4

@@ 4.3 Using more cameras
From the previous results, it can be seen that the number of cameras has a large impact on the performance of the algorithms. Especially mvor shows that using only three cameras while there are strong occlusions, leads to a large decrease in performance. From an application point of view, it is therefore recommended to use at least five or even more cameras, if this is possible.
To evaluate the impact of a high number of cameras, the egohumans [khirodkar2023ego] dataset was selected. It shows between 2-8 people in different indoor and outdoor settings (like building with lego bricks, fencing or playing tag, basketball, volleyball, badminton or tennis). They are watched by 8-20 cameras. The supervised space is up to $25\times 14\,m$ in size, which is much larger compared to the other datasets. The last recording of each setting was selected as test set, and each recording was subsampled into parts of $3s$ consecutive motions with larger gaps in between as well.
@@
Table 11: Transfer to egohumans [khirodkar2023ego]
Because the cameras use fisheye lenses, the images are distorted stronger than in the other datasets. Therefore a different undistortion method is required. Since this requires larger changes in the pipeline of most algorithms, only a subset of the better performing algorithms was tested on this dataset. In some cases the poses can be undistorted after the 2D estimation step, in other cases, the images need to be undistored instead. VoxelPose and Faster-VoxelPose, for which the images are undistorted beforehand, both have problems with a direct transfer to the legoassemble subset, but again show much better performance after synthetic training. The same holds true for VoxelPose on the tennis subset, Faster-VoxelPose on the other hand does not work at all here. The synthetic training always failed with an out-of-memory error, even at the lowest possible batch-size. For mvpose the number of views was too high, the six cameras from tsinghua already took up the complete GPU memory. TEMPO [choudhury2023tempo] described an evaluation on this dataset, in which they reported an average transfer error of 120mm120\,$m$ (or 36.7mm36.7\,$m$ after training) on this dataset, but they did not provide further information about the evaluation in the paper or their code. VoxelKeypointFusion proofs its generalization capabilities with a very good performance, which is only surpassed by RapidPoseTriangulation.
@@ 4.4 About real-time performance
Regarding the real-time performance with many cameras, the results show, that while the 3D part of RapidPoseTriangulation is very fast and at most takes 1ms1\,ms, the majority of the time is currently spent for 2D pose estimation.
In comparison to the original implementation of RTMPose the preprocessing pipeline and network interfacing are already optimized. Some parts of the preprocessing were moved to the GPU, and the image is transferred as uint8 type to decrease transfer time. The box cropping and padding were optimized, and for image resizing nearest-neighbor instead of linear interpolation is used for further speedups. Everything was implemented in C++ as well. Cropping and resizing directly on the GPU did not show notable speedups, it was counterweighted by an increased transfer time for the full image.
To further reduce the input latency in real applications, the Bayer format, which is the default layout of image sensors in most cameras, is used directly. Instead of having three RGB channels, this format only has one channel, in which every other pixel has a single red, green, or blue color only. The conversion into RGB is now integrated into the preprocessing pipeline. For the dataset evaluation, the RGB images were first converted back to Bayer encoding. Interestingly this back-and-forth conversion notably improved the MPJPE score in many datasets, presumably because the demosaicing algorithm of OpenCV is better than the default one of many cameras. The demosaicing time is already included in the FPS values of the previous tables. In the case of FullHD resolution, it is around 0.2ms0.2\,ms per image. With 4K inputs, like in the egohumans dataset, it increases to a notable 1.4ms1.4\,ms per image, so demosaicing after cropping and resizing, or directly using Bayer images as network input, might be future optimization possibilities.
@@
Another important latency factor in real-time systems is the transfer time of the image from the camera to the computer, especially if multiple cameras feed one host. Transferring a single FullHD image with $1920\times 1080\times 1$ pixels in Bayer encoding over a 1Gb/s1\,Gb/s network already takes around 17ms17\,ms. And this has to be stacked for every further camera. While compressing the image could reduce this time, it would on the other hand take extra processing time and reduce the image quality. A different solution could be adding a small computer to each camera, thus creating a smart-edge device, which estimates the poses for only one image and sends only a few keypoint coordinates to the main computer. Because only a single image is processed per time-step, this also would require less potent hardware, a Raspberry Pi with the AI-Hat could already be enough. If the price is not a concern, each camera could be equipped with one standard GPU though. The RTX-3090 used before takes about 1ms1\,ms per image and 1ms1\,ms per person to estimate their 2D pose. With an RTX-4080 this was even $45\%$ faster, and with an RTX-5090 (which has about $2.5\times$ more FLOPS than the RTX-3090), it should be possible to achieve full image-to-pose 3D prediction in only one millisecond.
@@ 5 Whole-body estimation
As already stated in VoxelKeypointFusion [voxkeyfuse], a significant benefit of algorithmic approaches over the learning-based ones, is their ability to handle input keypoints that were not part of the training dataset.
To evaluate the performance, the h3wb dataset [Zhu_2023_ICCV] was used, which extended a part of the human36m dataset with whole-body keypoints (17 body, 6 foot, 68 face, 42 hand). For simplification of the labeling process, only frames with good visibility of the actor were used, therefore the dataset is somewhat simpler than the original one (the default model of RapidPoseTriangulation reaches a MPJPE of 23.7mm23.7\,$m$, and the authors estimated around 17mm17\,$m$ labeling error). The whole-body results can be found in Table 12. Since this dataset is normally used for single-view 2D-to-3D pose lifting, no other comparable results were found.
One problem of VoxelKeypointFusion are the artifacts of the voxelization process, which due to the discretization pushes some keypoints closer together than they really are. This was especially visible at the finger keypoints, which were often clustered into agglomerations, instead of showing the clear lines of the fingers. RapidPoseTriangulation does not have this problem anymore, since it directly triangulates in continuous coordinates. On the other hand, smaller errors in the 2D keypoint prediction can lead to larger errors in 3D. For example if a keypoint is visible from three views, VoxelKeypointFusion merged them all at once using heatmap-beams, so if there are small 2D position errors the beams of all three views still partially overlap. RapidPoseTriangulation on the other hand triangulates each pair of views separately, and merges them afterward, but the errors do not necessarily cancel themselves out again. This is more severe if the keypoints can be only seen in very few views, like in this dataset, since it is not possible to detect and drop outliers before the merging step.
@@ 6 Conclusion
In a broad evaluation of the algorithm on several datasets it showed state-of-the-art performance. It scales well with an increased number of cameras and is especially suited for real-time applications. Besides that, the results also show that a relatively simple algebraic approach still can outperform many recent fully differentiable or end-to-end learned multi-view methods. This suggests that the current trend of increasing model complexity does not necessarily translate into better exploitation of the geometric structure.
@@ A.1 Number of cameras
For completeness an experiment with panoptic using $3$, 55, 77, $10$, or $31$ cameras, was executed. As expected, the results show, that more cameras lead to better results. The frame-rate also slows down, with the 2D-part having the most impact, and the 3D-part having a slight exponential increase, even though its impact is still very small.
For real-world applications this experiment is not very meaningful though, since the required number of cameras is highly dependent on the targeted environment. Depending on the setup, $3$ cameras can be enough or much too few, as in campus versus mvor. VP and F-VP both did an experiment for $3$ cameras on panoptic in which they had only very small performance drops, but their performance on campus and mvor showed a large drop, even though they also have $3$ cameras. For a real-world application, it therefore is better to first select a similar dataset, and then evaluate the impact of the number of cameras, and also their best positioning. Since it’s not possible to evaluate all combinations of datasets, cameras and positions, the recommendation for developers is instead to use the provided source-code and run a personalized evaluation on the most similar dataset.
@@ A.2 Algorithmic ablations
In general the algorithm is very robust against small changes, and only a few of them have a notable impact on the results. Ignoring the 2D confidence scores instead to integrating them into the 3D keypoint score, results in a slightly higher MPJPE. Removing the pair pre-filtering step, which removes pairs of poses before the triangulation, has minimal impact on most scores, but notably increases the triangulation time, since more pairs have to be calculated. The two outlier removal steps, before merging the proposals of one group to a single person notably improves the accuracy. Lowering the error threshold, that filters poor triangulations before merging, slightly increases MPJPE and the invalid count. The postprocessing steps of dropping clearly invalid persons or replacing clearly wrong joints by with their neighbors also improves the MPJPE and reduces the invalid person count in some datasets. Using only the triangulation error or only the reprojection error for the 3D keypoint score calculation, instead of a combination of both, has a slightly negative impact on the MPJPE. Using only the triangulation score is slightly faster though, because the more complex reprojection can be skipped. Especially in datasets with more cameras it is recommended to use increase the minimal required group size. This results in dropping of invalid 3D proposals which do not originate from the same person, but have a similar pose in two or more views. The higher the number of cameras the higher the chance for it. Since they are from different persons, they are triangulated to some mathematically valid positions somewhere in the 3D space. Proposals from the same person on the other hand start to build clusters around their real position, by which they can be recognized as such. The tracking step has only a minimal impact on most datasets, since it does not smooth the movement trajectories. At fast movements it might slightly reduce the position accuracy because it can slow down some position changes. On the other hand it can prevent physically impossible fast jitter movements. In case a person is not visible for a few frames, it can also prevent the person from being lost completely, since their last position is returned then, but it will also keep persons moving out of the room alive for a few extra frames. Therefore tracking is considered as an optional extension.

### v5

@@ 4.3 Using more cameras
The previous results demonstrate that the number of cameras has a large impact on the performance of the algorithms. Especially mvor shows that using only three cameras in the presence of strong occlusions leads to a large decrease in performance. From an application point of view, it is therefore recommended to use at least five or even more cameras, if this is possible.
To evaluate the impact of a high number of cameras, the egohumans [16] dataset was selected. It shows between 2-8 people in different indoor and outdoor settings (like building with lego bricks, fencing, or playing tag, basketball, volleyball, badminton, or tennis). They are watched by 8-20 cameras. The supervised space is up to $25\times 14\,m$ in size, which is much larger compared to the other datasets. The last recording of each setting was selected as test set, and each recording was subsampled into parts of $3s$ consecutive motions with larger gaps in between as well.
@@
Table 11: Transfer to egohumans [16]
Because the cameras use fisheye lenses, the images are more strongly distorted than in the other datasets. Therefore, a different undistortion method is required. Since this requires larger changes in the pipeline of most algorithms, only a subset of the better performing algorithms was tested on this dataset. In some cases the poses can be undistorted after the 2D estimation step, in other cases, the images need to be undistorted instead. VoxelPose and Faster-VoxelPose, for which the images are undistorted beforehand, both have problems with a direct transfer to the legoassemble subset, but again show much better performance after synthetic training. The same holds true for VoxelPose on the tennis subset, Faster-VoxelPose on the other hand does not work at all here. The synthetic training always failed with an out-of-memory error, even at the lowest possible batch-size. For mvpose the number of views was too high, the six cameras from tsinghua already took up the complete GPU memory. TEMPO [8] described an evaluation on this dataset, in which they reported an average transfer error of 120mm120\,$m$ (or 36.7mm36.7\,$m$ after training) on this dataset, but they did not provide further information about the evaluation in the paper or their code. VoxelKeypointFusion proves its generalization capabilities with a very good performance, which is only surpassed by RapidPoseTriangulation.
@@ 4.4 About real-time performance
Regarding real-time performance with many cameras, the results show that while the 3D part of RapidPoseTriangulation is very fast and takes at most 1ms1\,ms, the majority of the time is currently spent on 2D pose estimation.
In comparison to the original implementation of RTMPose, the preprocessing pipeline and network interfacing are already optimized. Some parts of the preprocessing were moved to the GPU, and the image is transferred as uint8 type to decrease transfer time. The box cropping and padding were optimized, and for image resizing nearest-neighbor instead of linear interpolation is used for further speedups. Everything was implemented in C++ as well. Cropping and resizing directly on the GPU did not show notable speedups, as transferring the full image counterweighted its benefits.
To further reduce the input latency in real applications, the Bayer format, which is the default layout of image sensors in most cameras, is used directly. Instead of having three RGB channels, this format only has one channel, in which every other pixel has a single red, green, or blue color only. The conversion into RGB is now integrated into the preprocessing pipeline. For the dataset evaluation, the RGB images were first converted back to Bayer encoding. Interestingly, this back-and-forth conversion notably improved the MPJPE score in many datasets, presumably because the demosaicing algorithm of OpenCV is better than the default one of many cameras. The demosaicing time is already included in the FPS values of the previous tables. In the case of FullHD resolution, it is around 0.2ms0.2\,ms per image. With 4K inputs, like in the egohumans dataset, it increases to a notable 1.4ms1.4\,ms per image, so demosaicing after cropping and resizing, or directly using Bayer images as network input, might be future optimization possibilities.
@@
Another important latency factor in real-time systems is the transfer time of the image from the camera to the computer, especially if multiple cameras feed one host. Transferring a single FullHD image with $1920\times 1080\times 1$ pixels in Bayer encoding over a 1Gb/s1\,Gb/s network already takes around 17ms17\,ms. And this has to be stacked for every further camera. While compressing the image could reduce this time, it would require extra processing time and reduce the image quality. A different solution could be adding a small computer to each camera, thus creating a smart-edge device, which estimates the poses for only one image and sends only a few keypoint coordinates to the main computer. Because only a single image is processed per time-step, this also would allow less powerful hardware. A Raspberry Pi with the AI-Hat could already be sufficient. If the price is not a concern, each camera could be equipped with one standard GPU though. The RTX-3090 used before takes about 1ms1\,ms per image and 1ms1\,ms per person to estimate their 2D pose. With an RTX-4080 this was even $45\%$ faster, and with an RTX-5090 (which has about $2.5\times$ more FLOPS than the RTX-3090), it should be possible to achieve full image-to-pose 3D prediction in only one millisecond.
@@ 5 Whole-body Estimation
As already stated in VoxelKeypointFusion [2], a significant benefit of algorithmic approaches over the learning-based ones, is their ability to handle input keypoints that were not part of the training dataset, such as whole-body joints.
To evaluate the performance, the h3wb dataset [31] was used, which extended a part of the human36m dataset with whole-body keypoints (17 body, 6 foot, 68 face, 42 hand). For simplification of the labeling process, only frames with good visibility of the actor were used. Therefore, the dataset is somewhat simpler than the original one (the default model of RapidPoseTriangulation reaches a MPJPE of 23.7mm23.7\,$m$, and the authors estimated around 17mm17\,$m$ labeling error). The whole-body results can be found in Table 12. Since this dataset is normally used for single-view 2D-to-3D pose lifting, no other comparable results were found.
One problem of VoxelKeypointFusion are the artifacts of the voxelization process, which, due to the discretization, push some keypoints closer together than they really are. This was especially visible at the finger keypoints, which were often agglomerated together instead of showing distinct finger lines. RapidPoseTriangulation does not have this problem anymore, since it directly triangulates in continuous coordinates. On the other hand, smaller errors in the 2D keypoint prediction can lead to larger errors in 3D. For example, if a keypoint is visible from three views, VoxelKeypointFusion merges them all at once using heatmap-beams. So if there are small 2D position errors, the beams of all three views still partially overlap. RapidPoseTriangulation on the other hand triangulates each pair of views separately, and merges them afterward, but the errors do not necessarily cancel themselves out again. This is more severe if the keypoints can only be seen in very few views, such as in this dataset, since it is not possible to detect and drop outliers before the merging step.
@@ 6 Conclusion
In a broad evaluation of the algorithm on several datasets, it showed state-of-the-art performance. It scales well with an increased number of cameras and is especially suited for real-time applications. Besides that, the results also show that a relatively simple algebraic approach still can outperform many recent fully differentiable or end-to-end learned multi-view methods. This suggests that the current trend of increasing model complexity does not necessarily translate into better exploitation of the geometric structure.
@@ A.1 Number of cameras
For completeness, an experiment with panoptic using $3$, 55, 77, $10$, or $31$ cameras was conducted. As expected, the results show that more cameras lead to better results. The frame-rate also slows down, with the 2D-part having the most impact, and the 3D-part having a slight exponential increase, even though its impact is still very small.
For real-world applications, this experiment is not very meaningful though, since the required number of cameras is highly dependent on the targeted environment. Depending on the setup, $3$ cameras can be enough or much too few, as in campus versus mvor. VP and F-VP both conducted an experiment for $3$ cameras on panoptic in which they had only very small performance drops, but their performance on campus and mvor showed a large drop, even though they also have $3$ cameras. For a real-world application, it is therefore better to first select a similar dataset and then evaluate the impact of the number of cameras and their best positioning. Since it is not possible to evaluate all combinations of datasets, cameras, and positions, the recommendation for developers is to use the provided source-code and run a personalized evaluation on the most similar dataset.
@@ A.2 Algorithmic ablations
In general, the algorithm is very robust against small changes, and only a few of them have a notable impact on the results. Ignoring the 2D confidence scores instead to integrating them into the 3D keypoint score, results in a slightly higher MPJPE. Removing the pair pre-filtering step, which removes pairs of poses before the triangulation, has minimal impact on most scores, but notably increases the triangulation time, since more pairs have to be calculated. The two outlier removal steps, before merging the proposals of one group to a single person, notably improve the accuracy. Lowering the error threshold, which filters poor triangulations before merging, slightly increases MPJPE and the invalid count. The postprocessing steps of dropping clearly invalid persons or replacing clearly wrong joints with their neighbors also improves the MPJPE and reduces the invalid person count in some datasets. Using only the triangulation error or only the reprojection error for the 3D keypoint score calculation, instead of a combination of both, has a slightly negative impact on the MPJPE. Using only the triangulation score is slightly faster though, because the more complex reprojection can be skipped. Especially in datasets with more cameras, it is recommended to increase the minimal required group size. This results in dropping invalid 3D proposals which do not originate from the same person, but have a similar pose in two or more views. The higher the number of cameras the higher the chance for it. Since they are from different persons, they are triangulated to some mathematically valid positions somewhere in the 3D space. Proposals from the same person on the other hand start to build clusters around their real position, by which they can be recognized as such. The tracking step has only a minimal impact on most datasets, since it does not smooth the movement trajectories. At fast movements it might slightly reduce the position accuracy because it can slow down some position changes. On the other hand it can prevent physically impossible fast jitter movements. In case a person is not visible for a few frames, it can also prevent the person from being lost completely, since their last position is returned then, but it will also keep persons moving out of the room alive for a few extra frames. Therefore tracking is considered as an optional extension.
Appendix B Further Remarks
The following section contains some further remarks about the algorithmic performance.
Method Dataset  PCP PCK@100/500  MPJPE Recall@100/500  Invalid  F1
mv3dpose shelf 92.2 (-4.9)    88.2 (-3.4) 98.4 (-4.8) 51.1 (-4.7)    90.6 (-4.2) 93.7 (-4.8) 40.6 (-3.7) 72.7 (+1.5)
PartAwarePose shelf 98.3 (0.0)    93.5 (+0.8) 99.0 (0.0) 48.1 (-3.3)    97.5 (-1.0) 99.2 (0.0) 49.4 (+2.0) 67.0 (-1.7)
RapidPoseTriangulation shelf 99.2    94.4 100 47.5    98.7 100 41.3 74.0
mv3dpose campus 87.1 (+3.0)    71.4 (+7.0) 91.4 (-2.0) 75.7 (-59)    81.6 (+19) 91.8 (-3.1) 17.5 (+7.4) 86.9 (-5.5)
PartAwarePose campus 94.2 (+1.0)    78.4 (0.0) 98.9 (0.0) 74.6 (-0.1)    92.0 (-1.9) 98.9 (0.0) 17.1 (-5.6) 90.2 (+3.4)
RapidPoseTriangulation campus 95.2    79.9 100 75.2    93.9 100 15.9 91.4
Table 15: Replacing the original 2D pose inputs with those of RTMPose, and the change to the original results.
B.1 Computation Comparison
The computation efficiency comparison was made as fair as technically possible. Results taken from [2] in Table 7 onwards used either RTX3090 (same as here), or GTX1080 (mvpose, mv3dpose, PartAwarePose) because they did not work on newer GPUs. In Table 2, the methods mvpose, VoxelPose, Faster-VoxelPose, PlaneSweepPose, VoxelKeypointFusion use a GPU, the rest is on CPU. The fastest three in the table ([4,5,26]) are closed source, so comparison on the same hardware is not possible. [4] used an Intel-i7 3.2GHz3.2\,GHz, with parallelization. [26] and PartAwarePose used a 3.7GHz3.7\,GHz CPU, [5] a 2.2GHz2.2\,GHz CPU, the AMD7900X used here has 4.7GHz4.7\,GHz, all those works used a single core. RapidPoseTriangulation is at least $30\times$ faster, which is much more than the $GHz$ difference, so the main speedup certainly is from the algorithm itself.
B.2 Different 2D Estimators
The referenced previous works use a wide variety of 2D detectors, and all results from Table 7 onwards are with the originally used detectors. VoxelPose, PRGnet, MvP, SelfPose3d all use ResNet [27], Faster-VP and TEMPO switched to the newer HRNet [22]. mvpose uses CPN [7], mv3dpose uses OpenPose [5], PartAwarePose uses HRNet [22]. VoxelKeypointFusion uses RTMPose [14], same as this work. In general, all previous works that switched their pose detector to newer models did not investigate this impact on performance. Speaking from experience, it is reasonable that this is not a standard practice in the field, because it takes a lot of effort to adjust often poorly documented code to inject different poses into it. For this reason, this work only conducted a small experiment with mv3dpose and PartAwarePose using exactly the same 2D poses as RapidPoseTriangulation, instead of the ones from their original OpenPose and HRNet pipelines (Table 15). For mv3dpose the results show improvements the localization accuracy (MPJPE), but a notable drop in person detection (Recall@500). The results of PartAwarePose show only marginal changes, mainly a slightly better MPJPE (mostly caused by better hips, while many joints even got worse). All scores are still clearly behind RapidPoseTriangulation, so the detector change is not the main reason for the better performance.
B.3 Additional Reasoning
The (voxel-based) learned methods all have generalization problems, they work well on the same dataset (panoptic), but show notable loss in setup switches. The synthetic training reduced this somewhat in many cases, but not to the full extend. They also do not allow using arbitrary keypoints, but require matching training datasets instead. This was already evaluated in more detail in [2]. And as VoxelKeypointFusion has shown, the problems do not come from the voxel representation itself. But even if those problems could be solved, those methods are rather slow (as shown in Table 2), which makes them less suitable for real-time applications, and, as VoxelKeypointFusion has shown, adding a lot more keypoints makes them even slower.
The algebraic methods often fail in person association. For example in Table 2+3 from [2] mv3dpose and PartAwarePose are not able to detect all persons (Recall@500), and as shown here in Table 15, this is not due to the 2D detector. The same problem can be seen in Table 5 from [2] or Table 9. The better performance of RapidPoseTriangulation in this regard can be explained by the different matching strategy. Unlike the others, RPT triangulates all possible view-pairs to 3D poses, and then drops those with large errors (steps 1-8). So if a person is seen in only two images it is very likely to be reconstructed. This makes RapidPoseTriangulation so reliable, which is why this approach was chosen. Grouping all those 3D proposals to persons (step 9) is very robust as well, because it is highly unlikely that two persons have very similar 3D poses while standing at basically the same spot. Because there are no projection caused similarities this is more reliable than 2D pose matching across views, especially in more crowded scenes, which have a higher probability of similar looking persons, and therefore more false matches. And compared to appearance based matching, it has no problems with similar looking persons (like in MVOR), and it is orders of magnitudes faster. While a larger number of persons also affects the likelihood of wrong 3D proposals generated in RapidPoseTriangulation, they can easily be filtered out by setting a higher minimal group size.
Regarding the joint accuracy (PCK@100), having multiple 3D proposals per person from the different view-pairs allows for a very efficient but robust outlier detection. At its core it’s based on majority voting, which removes very bad proposals, without them influencing the final position, which would not be possible with iterative matching or (weighted) averaging like in other works.